# Sensing atomic superfluid rotation beyond the standard quantum limit

### v1

@@ Appendix B Optimal homodyne detection angle for noise spectrum and sensitivity
Appendix C BAE steady state beyond RWA
@@
Comparing Eqs. (58) and (C), and comparing coefficients of $e^{i\delta t}$ up to the first order and neglecting fluctuations, the relevant relations are
@@
where the two values of $\Omega_{\rm eff}$ are for the cases with and and without inter-atomic interaction or collisions in the dynamics, respectively. Making the above substitutions, the steady-state populations $n_{\pm 1}=|a_{\pm 1}|^{2}$ under the above approximations, can be solved using a set of coupled cubic equations
@@
If the term $G$ is very small, the second term in the right hand side of Eqs. (72)-(73) vanishes and the steady state solution reduces to $\bar{a}=a_{s}=\dfrac{|\varepsilon_{\pm}|}{\delta^{2}+(\kappa/2)^{2}}$. This can then be considered to be the steady state value to linearize the optical operators in the BAE scheme. The validity of the above analytical expression can be checked by numerically solving the time-dependent set of differential equations in the mean field regime obtained from Eq. (57). The time-dependent complex solution $\alpha$ of the optical field are bounded in the phase space under the steady state oscillation amplitude $|a_{\pm}|$, as shown in Fig. 13. This shows that the solutions of the differential equations in the phase-space is strictly bounded by the steady-state amplitudes, which can then be used to find the operable monostable regions where rotational sensitivity is to be measured.
Appendix D Spectral density coefficients for the BAE scheme

### v2

@@ Appendix C Optimal homodyne detection angle for noise spectrum and sensitivity
Appendix D BAE steady state beyond RWA
@@
Comparing Eqs. (60) and (D), and comparing coefficients of $e^{i\delta t}$ up to the first order and neglecting fluctuations, the relevant relations are
@@
where the two values of $\Omega_{\rm eff}$ are for the cases with and and without inter-atomic interaction or collisions in the dynamics, respectively. Making the above substitutions, the steady-state populations $n_{\pm 1}=|\bar{a}_{\pm 1}|^{2}$ under the above approximations, can be solved using a set of coupled cubic equations
@@
If the term $G$ is very small, the second term in the right hand side of Eqs. (74)-(75) vanishes and the steady state solution reduces to $\bar{a}=a_{s}=\dfrac{|\varepsilon_{\pm}|}{\delta^{2}+(\kappa/2)^{2}}$. This can then be considered to be the steady state value to linearize the optical operators in the BAE scheme. The validity of the above analytical expression can be checked by numerically solving the time-dependent set of differential equations in the mean field regime obtained from Eq. (59). The time-dependent complex solution $\alpha$ of the optical field are bounded in the phase space under the steady state oscillation amplitude $|\bar{a}_{\pm}|$, as shown in Fig. 13. This shows that the solutions of the differential equations in the phase-space is strictly bounded by the steady-state amplitudes, which can then be used to find the operable monostable regions where rotational sensitivity is to be measured.
Appendix E Spectral density coefficients for the BAE scheme